OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) − 4 1 −1 −5 −25

OpenStudy (anonymous):

x=-1 do u need the steps

OpenStudy (precal):

distribute the 3 on the left side and distribute the -2 on the right side

OpenStudy (anonymous):

3 x+3 = -2 (x-1)-4

OpenStudy (anonymous):

3 x+3 = -2 x-2

OpenStudy (anonymous):

3 (1+x) = -4-2 (-1+x) 3+3 x = -4-2 (-1)-2 x 3+3 x = -4--2-2 x 3+3 x = -4+(-1)^(1+1) 2-2 x 3+3 x = -4+2 (-1)^(1+1)-2 x 3+3 x = -4+2-2 x 3+3 x = -4+2-2 x = 3 x+3 = -2 x-2